OpenStudy (anonymous):

I have a fill in the answer question for my study guide and I'd rather know how to figure this out now before the test. For f(x) = 1/x^2 - 3 find f(3) and f(2+h) I got f(3) I think, I just plugged in 3 for x but I'm confused how to get f(2 + h) it's a bit more confusing to me. Thank you for any help :)

OpenStudy (anonymous):

Did they give you any more information?

OpenStudy (anonymous):

No that's not that bad: \[f(x)=\frac{1}{x^2}\] So then, \(f(2+h)\) would be \(f(x)\) but putting \(2+h\) everywhere x is: \[f(2+h)=\frac{1}{(2+h)^2}=\frac{1}{h^2+4x+4}\]